# Cutting Sequence Diffuser: Sim-to-Real Transferable Planning for Object Shaping by Grinding

### v1

@@ IV-C Latent Encoding of Shape States
In grinding processes, shape representations such as point clouds and depth images used as states are high-dimensional. Directly using such high-dimensional states increases the training costs of a diffusion model. Therefore, a Variational Auto-Encoder (VAE) $\xi_{\psi,\mathrm{enc}}$ with parameters $\psi$ is used to compress state $\mathbf{s}_{t}$ into a latent feature $(\mathbf{z}_{t}=\xi_{\psi,\mathrm{enc}}(\mathbf{s}_{t}),\mathbf{z}_{t}\in\mathbb{R}^{dz})$. Where ${dz}$ is the dimension of the latent features. Thus, trajectory $\boldsymbol{\tau}$ used in a diffusion model is defined as $\boldsymbol{\tau}\coloneqq[\mathbf{z}_{t},\mathbf{a}_{t},\ldots,\mathbf{z}_{t+H-1},\mathbf{a}_{t+H-1}]$. In our experiments shown below, we employed point cloud $\mathcal{U}_{t}:=[{\mathbf{u}}_{m,t}]_{m=1}^{D}\subseteq\mathbb{R}^{3}$ as the state, where $D$ is the number of points and $\mathbf{u}_{m,t}$ is the $m$-th particle position.
@@ IV-D1 Cost Design
This section describes the types of cost functions used as guides for trajectory generation by a diffusion model. Total cost $c(\boldsymbol{\tau})$ for $t$ to $t+H-1$ is calculated by summing each cost: $c(\boldsymbol{\tau})=\lambda_{\mathrm{sm}}c_{\mathrm{sm}}(\boldsymbol{\tau})+\sum_{l=t}^{t+H-1}\sum_{k}\lambda_{k}c_{k}(\boldsymbol{\tau}[l])$, $k=\{\mathrm{state},\mathrm{col},\mathrm{vol}\}$. Where $\lambda$ are the weights of each cost.
@@
$d_{\text{len}}({a}_{l}^{j},{a}_{l+1}^{j})=\begin{cases}\|{a}_{l}^{{j}}-{a}_{l+1}^{j}\|_{2}^{2}&\text{if}~{}|{a}_{l}^{{j}}-{a}_{j+1}^{j}|>a^{j_{\text{max}}},\\ 0&\text{otherwise}.\end{cases}$ (12)
@@ IV-D2 Trajectory Generation with Two-step Guide
The proposed method increases task horizon $T$ by constraining the removal volume to reduce the reality gap. Therefore, as shown in §IV-A, the task horizon is divided into practical planning horizon lengths $H$. However, the grinding process is a goal-conditioned trajectory planning problem where only current shape $\mathbf{s}_{t}$ and final target shape $\mathbf{s}_{\mathrm{target}}$ can be used as state constraints. Thus, if we naively set the final target shape as the end of planning horizon constraint $t+H-1$, CSD tries to achieve object shaping, which normally requires $T$ steps, within a planning horizon $H$. This typically occurs when the planning horizon is shorter than the task horizon (i.e., $t+H-1\ll T$). Hence, such an approach can generate cutting sequences that remove a large volume at once, potentially causing significant removal resistance.
To address this issue, we execute trajectory planning with two-step guidance. Initially, at $t=0$, we set planning horizon $H$ equal to task horizon $T$ and generate a trajectory for the entire task horizon to obtain the reference shape states for each time step. Next, we plan the $H$-horizon trajectory by referring to the obtained shape as the state constraint at the end of the planning horizon, $t+H-1$. Thus, we use the shape state obtained by planning the trajectory with the entire task horizon as the state constraint at the end of the planning horizon. This approach allows for the suppression of cutting actions with large removal volumes at once. As a result, it may reduce removal resistance and promote smooth trajectory generation.
@@ V Robotic Rough-Grinding Environments
Shape A with ASA Shape A with PC Shape B with ASA Shape C with ASA
@@
Const-RS 1.42 ±$\pm 0.004$ 691.82 ±$\pm 9.03$ 1.42 ±$\pm 0.002$ 691.73 ±$\pm 8.61$ 1.89 ±$\pm 0.01$ 435.41 ±$\pm 1.95$ 1.60 ±$\pm 0.004$ 289.75 ±$\pm 5.04$
CSA-MBRL 0.96 ±$\pm 0.01$ 140.21 ±$\pm 1.16$ 1.07 ±$\pm 0.02$ 140.94 ±$\pm 0.98$ 1.52 ±$\pm 0.05$ 93.89 ±$\pm 0.88$ 1.40 ±$\pm 0.03$ 70.87 ±$\pm 0.94$
CSD (Ours) 0.94 ±$\pm 0.02$ 46.98 ±$\pm 0.81$ 0.95 ±$\pm 0.01$ 47.55 ±$\pm 0.91$ 1.31 ±$\pm 0.02$ 39.39 ±$\pm 0.81$ 1.18 ±$\pm 0.02$ 28.17 ±$\pm 0.64$
@@
Shape A with ASA Shape A with PC
@@
CSD w/ One-step Guide 0.074 ±$\pm 0.006$ 0.077 ±$\pm 0.011$
CSD w/ Two-step Guide 0.051 ±$\pm 0.006$ 0.044 ±$\pm 0.005$
@@ VI-C Data Collection and Action Planning Settings
We randomly sampled actions with $\epsilon_{\mathrm{vol}}=1.0$ and did not overcut the target shape of Shape A as a reference shape ${\mathbf{s}}_{\mathrm{ref}}$ by more than $d_{\mathrm{vol}}=0.3$. We prepared 9K trajectories with 251 steps per episode. Shape state ${\mathbf{s}_{t}}$ was compressed into latent features $(dz=64)$ by a VAE with a PointNet structure [24]. Following the same learning procedure as in [7, 8], we calculated the training loss of a 2D-Unet diffusion model using $c_{\mathrm{state}}$. Finally, we trained it in 2M steps with batch size $128$, $H=32$, and diffusion step $i=64$.
@@
Shape A with ASA Shape A with PC Shape B with ASA Shape C with ASA
@@
Const-RS 3.05 ±$\pm 0.15$ 1087.00 ±$\pm 4.20$ 2.58 ±$\pm 0.26$ 1090.55 ±$\pm 10.82$ 1.90 ±$\pm 0.19$ 900.00 ±$\pm 15.98$ 2.00 ±$\pm 0.10$ 662.28 ±$\pm 6.32$
CSA-MBRL 1.40 ±$\pm 0.09$ 89.13 ±$\pm 2.41$ 1.68 ±$\pm 0.16$ 87.48 ±$\pm 2.04$ 1.54 ±$\pm 0.02$ 90.33 ±$\pm 1.46$ 1.52 ±$\pm 0.20$ 68.02 ±$\pm 0.96$
CSD (Ours) 1.28 ±$\pm 0.06$ 49.61 ±$\pm 1.35$ 1.27 ±$\pm 0.03$ 45.87 ±$\pm 1.76$ 1.21 ±$\pm 0.02$ 35.46 ±$\pm 0.35$ 1.16 ±$\pm 0.02$ 28.63 ±$\pm 0.61$
@@
Shape A with ASA Shape A with PC
@@
CSD w/ One-step Guide 0.094 ±$\pm 0.012$ 0.086 ±$\pm 0.005$
CSD w/ Two-step Guide 0.046 ±$\pm 0.008$ 0.048 ±$\pm 0.006$

### v2

@@ IV-C Latent Encoding of Shape States
In grinding processes, shape representations such as point clouds and depth images used as states are high-dimensional. Directly using such high-dimensional states increases the training costs of a diffusion model. Therefore, a Variational Auto-Encoder (VAE) $\xi_{\psi,\mathrm{enc}}$ with parameters $\psi$ is used to compress state $\mathbf{s}_{t}$ into a latent feature $(\mathbf{z}_{t}=\xi_{\psi,\mathrm{enc}}(\mathbf{s}_{t}),\mathbf{z}_{t}\in\mathbb{R}^{dz})$. Where dz{dz} is the dimension of the latent features. Thus, trajectory $\boldsymbol{\tau}$ used in a diffusion model is defined as $\boldsymbol{\tau}\coloneqq[\mathbf{z}_{t},\mathbf{a}_{t},\ldots,\mathbf{z}_{t+H-1},\mathbf{a}_{t+H-1}]$. In our experiments shown below, we employed point cloud $\mathcal{U}_{t}:=[{\mathbf{u}}_{m,t}]_{m=1}^{D}\subseteq\mathbb{R}^{3}$ as the state, where $D$ is the number of points and $\mathbf{u}_{m,t}$ is the $m$-th particle position.
@@ IV-D1 Cost Design
This section describes the types of cost functions used as guides for trajectory generation by a diffusion model. Total cost $c(\boldsymbol{\tau})$ for $t$ to t+H1t+H-1 is calculated by summing each cost: $c(\boldsymbol{\tau})=\lambda_{\mathrm{sm}}c_{\mathrm{sm}}(\boldsymbol{\tau})+\sum_{l=t}^{t+H-1}\sum_{k}\lambda_{k}c_{k}(\boldsymbol{\tau}[l])$, $k=\{\mathrm{state},\mathrm{col},\mathrm{vol}\}$. Where $\lambda$ are the weights of each cost.
@@
$d_{\text{len}}({a}_{l}^{j},{a}_{l+1}^{j})=\begin{cases}\|{a}_{l}^{{j}}-{a}_{l+1}^{j}\|_{2}^{2}&\hskip-5.406pt\text{if}~|{a}_{l}^{{j}}-{a}_{j+1}^{j}|\hskip-1.13809pt>\hskip-1.13809pta^{j_{\text{max}}},\\ 0&\hskip-5.69054pt\text{otherwise}.\end{cases}$ (12)
@@ IV-D2 Trajectory Generation with Two-step Guide
The proposed method increases task horizon $T$ by constraining the removal volume to reduce the reality gap. Therefore, as shown in §IV-A, the task horizon is divided into practical planning horizon lengths $H$. However, the grinding process is a goal-conditioned trajectory planning problem where only current shape $\mathbf{s}_{t}$ and final target shape $\mathbf{s}_{\mathrm{target}}$ can be used as state constraints. Thus, if we naively set the final target shape as the end of planning horizon constraint t+H1t+H-1, CSD tries to achieve object shaping, which normally requires $T$ steps, within a planning horizon $H$. This typically occurs when the planning horizon is shorter than the task horizon (i.e., $t+H-1\ll T$). Hence, such an approach can generate cutting sequences that remove a large volume at once, potentially causing significant removal resistance.
To address this issue, we execute trajectory planning with two-step guidance. Initially, at $t=0$, we set planning horizon $H$ equal to task horizon $T$ and generate a trajectory for the entire task horizon to obtain the reference shape states for each time step. Next, we plan the $H$-horizon trajectory by referring to the obtained shape as the state constraint at the end of the planning horizon, t+H1t+H-1. Thus, we use the shape state obtained by planning the trajectory with the entire task horizon as the state constraint at the end of the planning horizon. This approach allows for the suppression of cutting actions with large removal volumes at once. As a result, it may reduce removal resistance and promote smooth trajectory generation.
@@ V Robotic Rough-Grinding Environments
Shape A with ASA
Shape A with PC
Shape B with ASA
Shape C with ASA
@@
Const-RS 1.42 $\pm 0.004$ 691.82 $\pm 9.03$ 1.42 $\pm 0.002$ 691.73 $\pm 8.61$ 1.89 $\pm 0.01$ 435.41 $\pm 1.95$ 1.60 $\pm 0.004$ 289.75 $\pm 5.04$
CSA-MBRL 0.96 $\pm 0.01$ 140.21 $\pm 1.16$ 1.07 $\pm 0.02$ 140.94 $\pm 0.98$ 1.52 $\pm 0.05$ 93.89 $\pm 0.88$ 1.40 $\pm 0.03$ 70.87 $\pm 0.94$
CSD (Ours) 0.94 $\pm 0.02$ 46.98 $\pm 0.81$ 0.95 $\pm 0.01$ 47.55 $\pm 0.91$ 1.31 $\pm 0.02$ 39.39 $\pm 0.81$ 1.18 $\pm 0.02$ 28.17 $\pm 0.64$
@@
Shape A with ASA
Shape A with PC
@@
CSD w/ One-step Guide 0.074 $\pm 0.006$ 0.077 $\pm 0.011$
CSD w/ Two-step Guide 0.051 $\pm 0.006$ 0.044 $\pm 0.005$
@@ VI-C Data Collection and Action Planning Settings
We randomly sampled actions with $\epsilon_{\mathrm{vol}}=1.0$ and did not overcut the target shape of Shape A as a reference shape ${\mathbf{s}}_{\mathrm{ref}}$ by more than $d_{\mathrm{vol}}=0.3$. We prepared 9K trajectories with 251 steps per episode. Shape state $\mathbf{s}_{t}$} was compressed into latent features $(dz=64)$ by a VAE with a PointNet structure [24]. Following the same learning procedure as in [7, 8], we calculated the training loss of a 1D-Unet diffusion model using $c_{\mathrm{state}}$. Finally, we trained it in 2M steps with batch size $128$, $H=32$, and diffusion step $i=64$.
@@
Shape A with ASA
Shape A with PC
Shape B with ASA
Shape C with ASA
@@
Const-RS 3.05 $\pm 0.15$ 1087.00 $\pm 4.20$ 2.58 $\pm 0.26$ 1090.55 $\pm 10.82$ 1.90 $\pm 0.19$ 900.00 $\pm 15.98$ 2.00 $\pm 0.10$ 662.28 $\pm 6.32$
CSA-MBRL 1.40 $\pm 0.09$ 89.13 $\pm 2.41$ 1.68 $\pm 0.16$ 87.48 $\pm 2.04$ 1.54 $\pm 0.02$ 90.33 $\pm 1.46$ 1.52 $\pm 0.20$ 68.02 $\pm 0.96$
CSD (Ours) 1.28 $\pm 0.06$ 49.61 $\pm 1.35$ 1.27 $\pm 0.03$ 45.87 $\pm 1.76$ 1.21 $\pm 0.02$ 35.46 $\pm 0.35$ 1.16 $\pm 0.02$ 28.63 $\pm 0.61$
@@
Shape A with ASA
Shape A with PC
@@
CSD w/ One-step Guide 0.094 $\pm 0.012$ 0.086 $\pm 0.005$
CSD w/ Two-step Guide 0.046 $\pm 0.008$ 0.048 $\pm 0.006$